Circular No. 7138

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V2134 OPHIUCHI = MXB 1659–29 J. in ’t Zand and J. Heise, Space Research Organization of the Nether-

lands (SRON), Utrecht; M. J. S. Smith, BeppoSAX Science Operations Center (SOC), Rome, and SRON; M. Cocchi and L. Natalucci, Istituto di Astrofisica Spaziale, CNR, Rome; and G. Celidonio, SOC, communicate: “During Apr. 2.06–3.47 UT, a transient x-ray source was detected through an observation of the Galactic-center field with Wide Field Camera (WFC) unit 1 on BeppoSAX. The location is α = 17^{h}02^{m}06^{s}, δ = −29^{o}56′.7 (equinox 2000.0; 99-percent confidence error radius 2′). This coincides with the position of MXB 1659–29, a transient x-ray burster last seen active 21 years ago. The average WFC-measured flux is 18 mCrab (2–9 keV). There is variability of a factor of 2 on a 5-hr timescale. Since its launch in 1996, WFC did not detect any other bursts from this position.”

T. Augusteijn, European Southern Observatory; L. Freyhammmer, Vrije Universiteit, Brussels; and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht; report: “Optical observations of V2134 Oph (the optical counterpart of the x-ray transient MXB 1659–29, detected by BeppoSAX), were made on Apr. 3.41 UT with the Dutch 0.90-m telescope at La Silla, showing the source to be bright at V = 18.3 ± 0.1. A spectrum (range 390–680 nm, resolution 0.6 nm) taken with the Danish 1.54-m telescope (+ DFOSC) on Apr. 4.39 showed the typical spectrum of a low-mass x-ray binary in an x-ray-bright phase, exhibiting emission lines of He II 468.6-nm, the C III/N III 464-/465-nm blend, and the Balmer series (Hα, with EW ∼ 0.5 nm and FWHM ∼ 700 km/s). Our observations confirm that the source is undergoing an outburst similar to the one detected in 1978 (Doxsey et al. 1979, Ap.J. 228, L67).”

SUPERNOVA 1999bh IN NGC 3435 G. Aldering and P. Nugent, Lawrence Berkeley National Laboratory

(LBNL); R. Ellis, Institute of Astronomy, University of Cambridge; and S. Perlmutter, LBNL, have obtained a spectrum (400–700 nm) of SN 1999bh (IAUC 7135) through service observations by D. Folha at the Issac Newton Telescope, La Palma, on Apr. 2 UT. The spectrum is that of a type-Ia supernova within one week after maximum light. The redshift of the supernova is ∼ 8400 km/s.

1999 April 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green